# Correspondence-Free, Function-Based Sim-to-Real Learning for Deformable Surface Control

### v1

@@ Abstract
This paper presents a correspondence-free, function-based sim-to-real learning method for controlling deformable freeform surfaces. Unlike traditional sim-to-real transfer methods that strongly rely on marker points with full correspondences, our approach simultaneously learns a deformation function space and a confidence map – both parameterized by a neural network – to map simulated shapes to their real-world counterparts. As a result, the sim-to-real learning can be conducted by input from either a 3D scanner as point clouds (without correspondences) or a motion capture system as marker points (tolerating missed markers). The resultant sim-to-real transfer can be seamlessly integrated into a neural network-based computational pipeline for inverse kinematics and shape control. We demonstrate the versatility and adaptability of our method on both vision devices and across four pneumatically actuated soft robots: a deformable membrane, a robotic mannequin, and two soft manipulators.
@@ I-A Kinematics for Shape Control
In the existing research works of soft robotics, high DoFs in the configuration space are typically computed by either reduced analytical models or discrete numerical models (see the comprehensive review in [11, 12]). Reduced analytical models are often adopted for soft manipulators with simple bending deformation (e.g., ref. [13, 7]), where the surface deformation is trivial in general. Differently, numerical simulation is usually employed to establish the mapping $\mathcal{S}(\mathbf{a})$ between the actuation parameters $\mathbf{a}$ and the deformed freeform shapes $\mathcal{S}$ (ref. [14, 15, 16, 17]). The differentiation of this forward kinematic mapping can then be used to solve the IK problem for determining the actuation parameters $\mathbf{a}$ according to the target shape $\mathcal{S}^{t}$. This was formulated as an optimization problem to iteratively run a simulator to determine the optimal actuation $\mathbf{a}_{opt}$ that minimizes the difference between $\mathcal{S}(\mathbf{a}_{opt})$ and $\mathcal{S}^{t}$ [18]. However, numerical simulation in general cannot be computed efficiently to realize a fast IK solver. A widely adopted strategy to mitigate the efficiency issue is to train a neural network (NN) by the simulation dataset as a surrogate model for the forward shape estimation (e.g., ref. [19, 20, 21]). This also adopts this strategy for IK computation.
@@ I-B Challenges of Sim-to-Real Transfer
Consequently, shape errors under large deformations are hard to be corrected effectively.
@@ I-C Our Method
Using a compact B-spline based representation of freeform surfaces, we propose a function-based pipeline that enables robust sim-to-real learning even when the input lacks of point correspondences (e.g., 3D scanned point clouds) or contains missing markers (e.g., MoCap data) / regions (e.g., 3D scan). First of all, it is based on a new architecture that learns the function space of shape deformation represented by radial basis functions (RBF). Specifically, each simulated free-form surface is approximated by a B-spline surface, the control points of which are employed as the shape descriptor in a lower dimension. The input of our sim-to-real network is the control points of a simulated body shape, and the output is the coefficients of RBFs that define a spatial warping function $\bm{\Phi}(\mathbf{p})$ by using the positions of kernel centers (Sec. II). The sim-to-real network $\mathcal{N}{s2r}$ consists of the components enclosed by the black dashed lines as shown in Fig. 2. Rather than relying on explicit point correspondences, it is trained jointly with a confidence estimation network $\mathcal{N}{conf}$ that predicts a confidence map. This map serves as a weighting function for the Chamfer distance between simulated and captured shapes. The training objective of sim-to-real transfer minimizes this weighted distance across numerous pairs of simulated and captured shapes, as illustrated in the differentiable alignment training module highlighted by the red dashed lines in Fig. 2. Details on the training process will be presented in Sec. III.
@@
After learning, the space warping function $\bm{\Phi}(\mathbf{p})$ is obtained in a closed-form and gives a continuous mapping from the shape of a simulated model to the shape of a corrected model that matches the reality. Working together with the forward kinematics prediction that is trained by simulation, this forms an end-to-end NN-based pipeline that can accurately predict the physical shapes of robotic deformable surfaces, which is crucial for efficient IK computation. This computation pipeline based on NN is analytically differentiable so that the gradient-based iteration for IK can be computed efficiently (details given in Appendix A-A and A-B). The quality of IK solutions has been verified by scanning the physical shape $\mathcal{S}^{p}$ to compare with the input target shape $\mathcal{S}^{t}$. All notations used in this paper are summarized in Table I.
@@
$\mathcal{N}_{fk}$ Forward kinematics network that predicts $\mathcal{S}^{c}$ from $\mathbf{a}$ $\mathcal{N}_{mk}$ The baseline network predicting positions of markers
$\bm{\gamma}=[\bm{\alpha}_{0},\ldots\bm{\beta}_{i}]$ Variables to determine the function space $\mathcal{N}_{rbf}$ Function prediction network that predicts $\bm{\gamma}$ by $\mathcal{S}^{c}$
$\mathbf{p}$, $\mathbf{p}^{*}$, $\mathbf{p}^{p}\in\mathbb{R}^{3}$ The surface point on $\mathcal{S}$, $\mathcal{S}^{*}$ and $\mathcal{S}^{p}$ respectively $\mathcal{N}_{s2r}$ Combination of $\mathcal{N}_{rbf}$ and $\bm{\Phi}(\cdot)$
$\bm{\Phi}(\cdot)$ The space warping function used to deform $\mathcal{S}$ to $\mathcal{S}^{*}$ $\{\mathbf{q}_{i}\}$ Centers of Gaussian kernels in $\bm{\Phi}(\mathbf{p})$
$(\mathbf{R},\mathbf{t})$ Rotation matrix and translation vector to re-pose $\mathcal{S}^{t}$ $\mathbf{c}^{*}$ Closet point of $\mathbf{p}^{*}$ on $\mathcal{S}^{t}$
$(u_{p},v_{p})$ A point in the parametric domain of B-spline surface $\mathbf{B}(\cdot)$ B-Spline function as decoder mapping $(u_{p},v_{p})$ to $\mathbf{p}$
@@ II Deformation Function Space for Sim-to-Real
This section introduces our function based sim-to-real learning for shape prediction. Given a set of points $\{\mathbf{p}\in\mathbb{R}^{3}\}$ on the free-form surface of a simulated model $\mathcal{S}$, we propose a method to learn a network $\mathcal{N}_{s2r}$ that can predict a deformation function mapping these points onto a corrected model $\mathcal{S}^{*}$ (i.e., the corrected shape that has eliminated the sim-to-real gap). The shape difference between $\mathcal{S}^{*}$ and the physical shape $\mathcal{S}^{p}$ should be minimized by the sim-to-real transfer. $\mathcal{N}_{s2r}$ is expected to be a general network [^1] Different networks still need be trained for different soft robotic setups. that can handle different shapes on the same hardware.
@@ II-A Compact Representation by B-Spline Surface
Using all sample points on the surface of $\mathcal{S}$ may introduce redundant information and therefore require a complex network for prediction. To overcome this issue, we first convert the freeform surface $\mathcal{S}$ of a soft deformable robot into a B-spline surface representation through 1) the parametrization taken at its rest shape (see Fig.3(a-c) for an example) and 2) the surface fitting conducted at every deformed shapes to determine control points of B-spline surface (denoted by $\mathcal{S}^{c}$ and see Fig.3(d)). The control points and B-Spline surfaces for four different soft robots are given in Fig.3(e). We then employ the control points of B-spline surface $\mathcal{S}^{c}$ as the shape descriptor of $\mathcal{S}$, with which any point in the parametric domain with the parameter $(u_{p},v_{p})$ can be mapped to a point $\mathbf{p}\in\mathbb{R}^{3}$ as $\mathbf{p}=\mathbf{B}(u_{p},v_{p},\mathcal{S}^{c})$. This function $\mathbf{B}(\cdot)$ is named as the B-spline based shape decoder in our computational pipeline (illustrated as the green block in Fig.2).
The dimension of $\mathcal{S}^{c}$ using B-spline representation is much lower than mesh representation. For instance, we use $30\times 30$ control points for all robots as shown in Fig.3(e). Note that, since the surfaces obtained via simulation retain an unchanged mesh topology (i.e., preserving correspondence), the computation of B-spline surface fitting is based on the parameterization determined by the rest shape of $\mathcal{S}$.
@@ II-B RBF-based Spatial Warping
Considering the gap between the simulated shape $\mathcal{S}$ and the physical shape $\mathcal{S}^{p}$, we propose to model it as a continuous spatial warping by the radial basis functions (RBF) as follows.
@@
with $N$ being the number of kernels employed for the space warping and $A=[\bm{\alpha}_{1},\bm{\alpha}_{2},\bm{\alpha}_{3}]$. The centers of Gaussian kernels, $\{\mathbf{q}_{i}\}$, are sparsely selected to cover the entire surface of a simulated model. The value of coefficient $c$ controlling the width of the Gaussian kernel is chosen as $3.0\times 10^{-5}$ by experiments. We remark that $\bm{\gamma}=[\bm{\alpha}_{0},\bm{\alpha}_{1},...\bm{\beta}_{i}]$ are variables of the function space, which can be determined by solving a linear system [26]. As we represent the free-form surface of each simulated model by a B-spline surface (Sec. II-A), the same kernel on different simulated shapes will have the same $(u,v)$-parameters but different center-positions in $\mathbb{R}^{3}$.
@@
[…]
  • This is not an end-to-end approach as it needs to solve a linear equation system when establishing a warping function for each simulated shape;
  • Training the marker prediction network $\mathcal{N}_{mk}$ needs sets of fully captured markers, which can be challenging on largely deformed models as discussed above.
A function space learning-based method is introduced below to solve these issues. Specifically, a network $\mathcal{N}_{rbf}$ is introduced to predict the value of the function variable $\bm{\gamma}$ according to the simulated shape $\mathcal{S}$.
@@ II-C NN-based Deformation Function Space
Our function-prediction based sim-to-real pipeline is formed by three building blocks: 1) the network $\mathcal{N}_{rbf}$, 2) the B-spline based shape decoder and 3) the RBF-based space warping (i.e., Eq.(1)). The diagram has been shown in the top of Fig.2. When a deformable robot is actuated into various shapes $\mathcal{S}_{j}$} by corresponding actuation $\{\mathbf{a}_{j}\}$, the function-prediction network $\mathcal{N}_{rbf}$ is trained on pairs of shape differences between the shape $\mathcal{S}_{j}^{*}$ predicted by the sim-to-real network $\mathcal{N}_{s2r}$ and the physically captured shape $\mathcal{S}_{j}^{p}$. More specifically, for the $j$-th frame, the surface of the simulated shape $\mathcal{S}_{j}$ is sampled as a set of points $\{(u_{i},v_{i})\}$ in its parametric domain. After applying sim-to-real correction, the resultant surface $\mathcal{S}_{j}^{*}$ is represented as a point set $\{\mathbf{p}^{*}_{i}\}$, where each point is given by $\mathbf{p}^{*}_{i}=\bm{\Phi}_{j}(\mathbf{B}(u_{i},v_{i},\mathcal{S}^{c}_{j}))$. The warping function $\bm{\Phi}_{j}$ is parameterized by $\bm{\gamma}$, which is predicted by the network as $\bm{\gamma}=\mathcal{N}_{rbf}(\mathcal{S}^{c}_{j})$. The corresponding physically acquired shape $\mathcal{S}_{j}^{p}$ is also represented as a point set $\{\mathbf{p}^{p}_{i}\}$. Details on how the shape difference between $\mathcal{S}_{j}^{*}$ and $\mathcal{S}_{j}^{p}$ is effectively evaluated for training $\mathcal{N}_{rbf}$ will be presented in Sec. III.
Given a well trained network $\mathcal{N}_{rbf}(\cdot)$, the sim-to-real transfer can be evaluated by the pipeline of Fig.2 in the inference phase as a network $\mathcal{N}_{s2r}(\cdot)$ that gives
@@
$(u_{p},v_{p})$ denotes a point sampled in the parametric domain of the free-form surface. $\{\mathbf{B}(u_{q},v_{q},\mathcal{S}^{c}(\mathbf{a})\}$ gives the set of kernel centers for the RBF-based warping function, which are defined by their parameters $(u_{q},v_{q})$. All operators in this sim-to-real pipeline is differentiable w.r.t. the actuation parameter $\mathbf{a}$. Detail analysis of differentiability can be found in Appendix A-B. In Sec.IV, we will present the detail of NN-based forward kinematics and the IK computing for shape control, which is based on minimizing the distance between $\mathbf{p}^{*}$ and its closest point $\mathbf{c}^{*}$ on the target shape $\mathcal{S}^{t}$.

### v2

@@ Abstract
This paper presents a correspondence-free, function-based sim-to-real learning method for controlling deformable freeform surfaces. Unlike traditional sim-to-real transfer methods that strongly rely on marker points with full correspondences, our approach simultaneously learns a deformation function space and a confidence map – both parameterized by a neural network – to map simulated shapes to their real-world counterparts. As a result, the sim-to-real learning can be conducted by input from either a 3D scanner as point clouds (without correspondences) or a motion capture system as marker points (tolerating missed markers). The resultant sim-to-real transfer can be seamlessly integrated into a neural network-based computational pipeline for inverse kinematics and shape control. We demonstrate the versatility and adaptability of our method on two vision devices and across four pneumatically actuated soft robots: a deformable membrane, a robotic mannequin, and two soft manipulators.
@@ I-A Kinematics for Shape Control
In the existing research works of soft robotics, high DoFs in the configuration space are typically computed by either reduced analytical models or discrete numerical models (see the comprehensive review in [11, 12]). Reduced analytical models are often adopted for soft manipulators with simple bending deformation (e.g., ref. [13, 7]), where the surface deformation is trivial in general. Differently, numerical simulation is usually employed to establish the mapping $\mathcal{S}(\mathbf{a})$ between the actuation parameters $\mathbf{a}$ and the deformed freeform shapes $\mathcal{S}$ (ref. [14, 15, 16, 17]). The differentiation of this forward kinematic mapping can then be used to solve the IK problem for determining the actuation parameters $\mathbf{a}$ according to the target shape $\mathcal{S}^{t}$. This was formulated as an optimization problem to iteratively run a simulator to determine the optimal actuation $\mathbf{a}_{opt}$ that minimizes the difference between $\mathcal{S}(\mathbf{a})$ and $\mathcal{S}^{t}$ [18]. However, numerical simulation in general cannot be computed efficiently to realize a fast IK solver. A widely adopted strategy to mitigate the efficiency issue is to train a neural network (NN) by the simulation dataset as a surrogate model for the forward shape estimation (e.g., ref. [19, 20, 21]). This paper also adopts this strategy for IK computation.
@@ I-B Challenges of Sim-to-Real Transfer
Consequently, a correspondence-free method for sim-to-real learning is needed to correct shape errors under large deformations.
@@ I-C Our Method
The focus of this paper is a new learning-based sim-to-real transfer method that corrects the shape errors produced by NN-based forward prediction trained on simulation data, thereby enabling efficient and accurate NN-based IK computation for deformation control of soft robots, rather than proposing a new control strategy.
Using a compact B-spline based representation of freeform surfaces, we propose a function-based pipeline that enables robust sim-to-real learning even when the input lacks of point correspondences (e.g., 3D scanned point clouds) or contains missing markers (e.g., MoCap data) / regions (e.g., 3D scan). First of all, it is based on a new architecture that learns the function space of shape deformation represented by radial basis functions (RBF). Specifically, each simulated free-form surface is approximated by a B-spline surface, the control points of which are employed as the shape descriptor in a lower dimension. The input of our sim-to-real network is the control points of a simulated body shape, and the output is the coefficients of RBFs that define a spatial warping function $\bm{\Phi}(\mathbf{p})$ by using the positions of kernel centers (Sec. II). The sim-to-real network $\mathcal{N}{s2r}$ consists of two components enclosed by the black dashed lines as shown in Fig. 2. Rather than relying on explicit point correspondences, it is trained jointly with a network $\mathcal{N}{conf}$ that predicts a confidence map. This map serves as a weighting function for evaluating the Chamfer distance between simulated and captured shapes. The training objective of sim-to-real transfer minimizes this weighted distance across numerous pairs of simulated and captured shapes, as illustrated in the differentiable alignment training module highlighted by the red dashed lines in Fig. 2. Details of this training process will be presented in Sec. III.
@@
After learning, the space warping function $\bm{\Phi}(\mathbf{p})$ is obtained in a closed-form and gives a continuous mapping from the shape of a simulated model to the shape of a corrected model that matches the reality. Working together with the forward kinematics prediction that is trained by simulation, this forms an end-to-end NN-based pipeline that can accurately predict the physical shapes of robotic deformable surfaces, which is crucial for efficient IK computation. This computation pipeline based on NN is analytically differentiable so that the gradient-based iteration for IK can be computed efficiently (details given in Sec. IV-A). The quality of IK solutions will be verified by scanning the physical shape $\mathcal{S}^{p}$ to compare with the input target shape $\mathcal{S}^{t}$. All notations used in this paper are summarized in Table I.
@@
$\mathcal{N}_{fk}$ Forward kinematics network that predicts $\mathcal{S}^{c}$ from $\mathbf{a}$ $\mathcal{N}_{rbf}$ Function prediction network that predicts $\bm{\gamma}$ by $\mathcal{S}^{c}$
$\bm{\gamma}=[\bm{\alpha}_{0},\ldots\bm{\beta}_{i}]$ Variables to determine the function space $\mathcal{N}_{s2r}$ Combination of $\mathcal{N}_{rbf}$ and $\bm{\Phi}(\cdot)$
$\mathbf{p}$, $\mathbf{p}^{*}$, $\mathbf{p}^{p}\in\mathbb{R}^{3}$ The surface point on $\mathcal{S}$, $\mathcal{S}^{*}$ and $\mathcal{S}^{p}$ respectively $\{\mathbf{q}_{i}\}$ Centers of Gaussian kernels in $\bm{\Phi}(\mathbf{p})$
$\bm{\Phi}(\cdot)$ The space warping function used to deform $\mathcal{S}$ to $\mathcal{S}^{*}$ $\mathbf{c}^{*}$ Closet point of $\mathbf{p}^{*}$ on $\mathcal{S}^{t}$
$(\mathbf{R},\mathbf{t})$ Rotation matrix and translation vector to re-pose $\mathcal{S}^{t}$ $\mathbf{B}(\cdot)$ B-Spline function as decoder mapping $(u_{p},v_{p})$ to $\mathbf{p}$
$(u_{p},v_{p})$ A point in the parametric domain of B-spline surface $M_{u},M_{v}$ # of control points along $u$ and $v$ directions respectively
@@ II Deformation Function Space for Sim-to-Real
This section introduces our function based sim-to-real learning for shape prediction. Given a set of points $\{\mathbf{p}\in\mathbb{R}^{3}\}$ on the free-form surface of a simulated model $\mathcal{S}$, we propose a method to learn a network $\mathcal{N}_{s2r}$ that can predict a deformation function for mapping these points onto a corrected model $\mathcal{S}^{*}$ (i.e., the corrected shape that has eliminated the sim-to-real gap). The shape difference between $\mathcal{S}^{*}$ and the physical shape $\mathcal{S}^{p}$ should be minimized by the sim-to-real transfer. $\mathcal{N}_{s2r}$ is expected to be a general network that can handle different shapes on the same hardware [^1] Different networks still need be trained for different soft robotic setups..
@@ II-A Compact Representation by B-Spline Surface
Using all sample points on the surface of $\mathcal{S}$ may introduce redundant information and therefore requires a complex network for prediction. To overcome this issue, we first convert the freeform surface $\mathcal{S}$ of a soft deformable robot into a B-spline surface representation through 1) the parametrization taken at its rest shape (see Fig.3(a-c) for an example) and 2) the surface fitting conducted at every deformed shapes to determine control points of B-spline surface (denoted by $\mathcal{S}^{c}$ and see Fig.3(d)). The control points and B-Spline surfaces for four different soft robots are given in Fig.3(e). We then employ the control points of B-spline surface $\mathcal{S}^{c}$ as the shape descriptor of $\mathcal{S}$, with which any point in the parametric domain with the parameter $(u_{p},v_{p})$ can be mapped to a point $\mathbf{p}\in\mathbb{R}^{3}$ as $\mathbf{p}=\mathbf{B}(u_{p},v_{p},\mathcal{S}^{c})$. According to the definition of a B-spline surface [26], we have
$\mathbf{p}=\mathbf{B}(u,v,\mathcal{S}^{c})=\sum_{i=1}^{M_{u}}\sum_{j=1}^{M_{v}}N_{i,k}(u)N_{j,l}(v)\mathcal{S}^{c}_{ij}.$ (1)
The sample point $\mathbf{p}$ on the B-spline surface is the sum of the linear combination of the B-spline basis function (i.e., $N_{ik}(u)$ and $N_{jl}(v)$) and the control point $\mathcal{S}^{c}_{ij}$. $k$ and $l$ denote the order of B-spline basis function in $u$ and $v$ direction, where cubic B-spline is chosen in the paper (i.e., $k=l=4$). $M_{u}$ and $M_{v}$ are the number of control points along $u$ and $v$ direction respectively. This function $\mathbf{B}(\cdot)$ is named as the B-spline based shape decoder in our computational pipeline (illustrated as the light green block at the top of Fig.2).
The dimension of $\mathcal{S}^{c}$ using B-spline representation as $M_{u}\times M_{v}$ is much lower than mesh representation. For instance, we use $30\times 30$ control points for all robots as shown in Fig.3(e). Note that, since the surfaces obtained via simulation retain an unchanged mesh topology (i.e., preserving correspondence), the computation of B-spline surface fitting is based on the parameterization determined by the rest shape of $\mathcal{S}$.
@@ II-B RBF-based Spatial Warping
Considering the gap between the simulated shape $\mathcal{S}$ and the physical shape $\mathcal{S}^{p}$, we propose to model it as continuous spatial warping by the radial basis functions (RBF) for any 3D point $\mathbf{p}\in\mathbb{R}^{3}$ as follows.
@@
with $N$ being the number of kernels employed for the space warping and $A=[\bm{\alpha}_{1},\bm{\alpha}_{2},\bm{\alpha}_{3}]$. The centers of Gaussian kernels, $\{\mathbf{q}_{i}=\mathbf{B}(u_{i},v_{i},\mathcal{S}^{c})\}$, are chosen by sparsely sampling across the UV domain to achieve comprehensive coverage of the simulated model’s surface with the control points $\mathcal{S}^{c}$. The number of kernels for different robots will be provided in Table II. The value of coefficient $c$ controlling the width of the Gaussian kernel is chosen as $3.0\times 10^{-5}$ by experiments. We remark that $\bm{\gamma}=[\bm{\alpha}_{0},\bm{\alpha}_{1},...\bm{\beta}_{i}]$ with $\bm{\alpha}_{j},\bm{\beta}_{i}\in\mathbb{R}^{3}$ are variables of the function space to be determined via the learning process. Note that the same kernel on different simulated shapes will have the same UV parameters but different center-positions in $\mathbb{R}^{3}$.
@@
With a determined warping function $\bm{\Phi}(\mathbf{p})$, we are able to map any point on the simulated surface $\mathcal{S}$ denoted as $\mathbf{p}\in\mathcal{S}$ to its corrected position as $\mathbf{p}^{*}\in\mathcal{S}^{*}$ – see Fig.4 for the illustration. However, different warping functions – by changing the values of $\gamma$ need to be employed for different simulated shapes. Specifically, we introduce a network $\mathcal{N}_{rbf}$ to predict the function coefficients $\bm{\gamma}$ according to the simulated shape $\mathcal{S}$ as illustrated in Fig.2.
@@ II-C NN-based Deformation Function Space
Our function-prediction based sim-to-real pipeline is formed by three building blocks: 1) the network $\mathcal{N}_{rbf}$, 2) the B-spline based shape decoder $\mathbf{B}(\cdot)$ and 3) the RBF-based space warping $\Phi(\cdot)$ (i.e., Eq.(2)). The diagram has been shown in the top of Fig.2. The purpose of network $\mathcal{N}_{rbf}$ is to generate shape-dependent RBF functions by determining their weights as
$\bm{\gamma}=\mathcal{N}_{rbf}(\mathcal{S}^{c})$ (3)
with its input as $M_{u}\times M_{v}$ control points of a simulated shape (denoted as $\mathcal{S}^{c}$) and the output $\bm{\gamma}=[\bm{\alpha}_{0},\bm{\alpha}_{1},...\bm{\beta}_{i}]$ as the collection of $N+4$ coefficient vectors in $\mathbb{R}^{3}$.
When a deformable robot is actuated into various shapes $\{\mathcal{S}_{j}\}$ by corresponding actuation $\{\mathbf{a}_{j}\}$, the function-prediction network $\mathcal{N}_{rbf}$ can be trained on pairs of shape differences between the shape $\mathcal{S}_{j}^{*}$ predicted by the sim-to-real network $\mathcal{N}_{s2r}$ and the physically captured shape $\mathcal{S}_{j}^{p}$. More specifically, for the $j$-th deformation, the surface of the simulated shape $\mathcal{S}_{j}$ is sampled as a set of points $\{(u_{i},v_{i})\}$ in its parametric domain. After applying sim-to-real correction, the resultant surface $\mathcal{S}_{j}^{*}$ is represented as a point set $\mathcal{S}_{j}^{*}=\{\mathbf{p}^{*}_{i}\}$, where each point can be computed by
$\color[rgb]{0,0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.0}\mathbf{p}_{i}^{*}=\Phi_{\bm{\gamma}}(\mathbf{p})=\Phi_{\mathcal{N}_{rbf}(\mathcal{S}_{j}^{c})}\left(\mathbf{B}(u_{i},v_{i},\mathcal{S}_{j}^{c})\right).$}} (4)
The warping function $\bm{\Phi}_{j}$ is parameterized by $\bm{\gamma}$, which is predicted by the network as $\bm{\gamma}=\mathcal{N}_{rbf}(\mathcal{S}^{c}_{j})$. The corresponding physically acquired shape $\mathcal{S}_{j}^{p}$ is also represented as a point set $\{\mathbf{p}^{p}_{k}\}$. For training the network $\mathcal{N}_{rbf}$, the shape differences between all pairs of point clouds as $S_{j}^{*}=\{\mathbf{p}_{k}^{*}\}$ and $S_{j}^{p}=\{\mathbf{p}^{p}_{k}\}$ are employed to define loss functions to be minimized. Note that our loss functions are defined in a way that does not require pre-defined correspondences between two point clouds and also tolerate missed regions. Details will be presented in Sec. III.
Given a well trained network $\mathcal{N}_{rbf}(\cdot)$, the sim-to-real transfer can be evaluated by the pipeline at the top of Fig.2 in the inference phase. For any point $(u_{p},v_{p})$ sampled in the UV domain on a simulated shape compactly encoded as a set of control points $\mathcal{S}^{c}$, we can obtain its predicted position by Eq.(4). By modeling $\mathcal{S}^{c}$ as a neural network–based function of the actuation vector $\mathbf{a}$, the above sim-to-real transfer process can be interpreted as a network $\mathcal{N}_{s2r}(\cdot)$ that gives
@@
All operators in this sim-to-real pipeline is differentiable w.r.t. the actuation parameter $\mathbf{a}$. Detail analysis of differentiability can be found in Appendix A-A. In Sec. IV, we will present the detail of NN-based forward kinematics and the IK computing for shape control, which is based on minimizing the distance between $\mathbf{p}^{*}$ and its closest point $\mathbf{c}^{*}$ on the target shape $\mathcal{S}^{t}$.